### v3

# Quantum HodgeRank: Topology-Based Rank Aggregation on Quantum Computers

Abstract
HodgeRank generalizes ranking algorithms, e.g. Google’s PageRank, to rank alternatives based on real-world (often incomplete) data using graphs and discrete exterior calculus. It analyzes multipartite interactions on high-dimensional networks with a complexity that scales exponentially with dimension. We develop a quantum algorithm that approximates the HodgeRank solution with complexity independent of dimension. Our algorithm extracts relevant information from the state such as the ranking consistency, which achieves a superpolynomial speedup over similar classical methods.
Introduction.—Ranking is crucial for informed decision-making and comparing alternatives. Google’s PageRank [1, 2] and Netflix’s recommendation system [3] are both, at their core, ranking algorithms. Ranking systems typically operate on noisy or incomplete data. For instance, consider ranking football teams based on match results: the data compares the teams pairwise, but some pairs may be missing, and thus the data is incomplete. Such ranking data appears commonly in systems with many alternatives, as many pairs might not have been explicitly compared.
HodgeRank [4] overcomes such limitations by utilizing discrete exterior calculus [5, 6] to aggregate incomplete data – e.g., the pairwise match results – to derive a global ranking. It also provides a consistency measure for its ranking, indicating whether the data implies a strong consensus ranking of alternatives or whether the data contains ranking inconsistencies, such as Condorcet paradoxes (rock-paper-scissors scenarios). HodgeRank generalizes PageRank and has applications in recommendation systems, social choice theory, arbitrage-free market detection, sports tournament ranking, online video quality assessments, market analysis, and protein folding analysis [4, 7, 8, 9, 10, 11].
The above examples involve ranking based on pairwise comparisons, but the HodgeRank tools of discrete exterior calculus on simplicial complexes generalize to higher-order networks [12, 13, 14, 15, 16]. These are used to model complex phenomena such as biological networks [17, 18] and chaotic Hamiltonian dynamics [19, 20], with further applications to machine learning [21, 22, 23]. However, higher-dimensional HodgeRank ($k$-HodgeRank, for network dimension $k$) computations are expensive: the complexity scales exponentially in $k$. This hinders practical applications and exploration of higher-order network models despite strong evidence for their importance in real-world systems [24, 25, 26].
In this Letter, we propose a quantum algorithm that prepares a quantum state proportional to the output of $k$-HodgeRank that is exponentially faster in $k$ than generalizations of classical methods. This advantage is maintained when extracting answers to specific questions, modulo some caveats discussed later. We give algorithms to (i) measure rankability metrics of the data, (ii) relatively rank a subset of alternatives, and (iii) return top-ranking alternatives. In particular, the first applicationis superpolynomially faster for large $k$, and polynomially faster for $k=1$ (corresponding to pairwise preference data), than comparable classical methods. This is similar to other quantum algorithms that estimate the fit quality of data [27, 28]. The latter applications are more heuristic and data-dependent. While we focus on ranking algorithms here, the tools developed in the Letter will have broader applications to higher-dimensional discrete exterior calculus, such as solving general least squares [29] and boundary value problems [30, 31].
HodgeRank.— We briefly describe HodgeRank and the associated mathematical problems (see [4] for details). HodgeRank outputs a global ranking of alternatives $\mathcal{V}:=\{v_{1},\dots,v_{n}\}$. The input to HodgeRank is aggregated pairwise preferences of voters, modelled by an edge flow $\bm{\mathrm{s}}^{1}:\mathcal{E}\to\mathcal{R}$ on a comparison graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. This corresponds to a vector $\bm{\mathrm{s}}^{1}\in\mathbb{R^{|\mathcal{E}|}}$ by (arbitrarily) orienting each edge in $\mathcal{E}$. The graph $\mathcal{G}$ need not be complete: not all pairs of alternatives must be compared directly. The value $\bm{\mathrm{s}}^{1}([v_{i},v_{j}])$ represents the preference for alternative $i$ over $j$, where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes edge orientation.
Given a comparison graph $\mathcal{G}$ and preference data $\bm{\mathrm{s}}^{1}$, HodgeRank solves a weighted $\ell_{2}$ minimization problem with solution
$\bm{\mathrm{s}}^{1}_{\mathrm{G}}=\operatorname*{arg\,min}_{\bm{\mathrm{x}}\in\mathrm{im}(\bm{\mathrm{B}}_{1}^{\dagger})}\left[\sum_{1\leq i<j\leq n}w_{ij}\left(x_{ij}-\bm{\mathrm{s}}^{1}([v_{i},v_{j}])\right)^{2}\right],$ (1)
where $w_{ij}$ represents the weight associated to the data on (oriented) edge $[v_{i},v_{j}]$ and $\bm{\mathrm{B}}_{k}$ is a matrix representation of the $k$-th boundary operator of a clique complex $\mathcal{K}_{n}(\mathcal{G})$ (see Supplemental Material [32]). The space $\mathrm{im}(\bm{\mathrm{B}}_{1}^{\dagger})$ contains so-called conservative edge flows, which means there exists $\bm{\mathrm{s}}^{0}_{*}\in\mathbb{R}^{|\mathcal{V}|}$ such that $\bm{\mathrm{s}}^{1}_{\mathrm{G}}([v_{i},v_{j}])=\bm{\mathrm{s}}^{0}_{*}(v_{i})-\bm{\mathrm{s}}^{0}_{*}(v_{j})$. In a discrete exterior calculus sense, $\bm{\mathrm{s}}^{1}_{\mathrm{G}}$ is an analogue of the gradient of the potential function $\bm{\mathrm{s}}^{0}_{*}$.
HodgeRank then outputs the potential (or score) function $\bm{\mathrm{s}}^{0}_{*}$ from Eq. (1) by $\bm{\mathrm{B}}_{1}^{\dagger}\bm{\mathrm{s}}^{0}_{\ast}=\bm{\mathrm{s}}^{1}_{\mathrm{G}}$. This defines a global ranking of the alternatives: the score of alternative $v_{i}$ is $\bm{\mathrm{s}}^{0}_{*}(v_{i})$, and $v_{i}$ outranks $v_{j}$ whenever $\bm{\mathrm{s}}^{0}_{*}(v_{i})\geq\bm{\mathrm{s}}^{0}_{*}(v_{j})$. In this Letter, we set $w_{ij}=0$ if $v_{i},v_{j}$ are not compared, and 1 otherwise; we discuss this assumption in Supplemental Material [32]. Under this assumption, $\bm{\mathrm{s}}^{0}_{\ast}=\big{(}\bm{\mathrm{B}}_{1}\bm{\mathrm{B}}_{1}^{\dagger}\big{)}^{+}\bm{\mathrm{B}}_{1}\bm{\mathrm{s}}^{1}$, where $+$ denotes the Moore-Penrose pseudoinverse. The relative length of $\bm{\mathrm{s}}^{1}_{\mathrm{G}}$ compared to $\bm{\mathrm{s}}^{1}$ measure the rankability of the alternatives according to $\bm{\mathrm{s}}^{1}$.
In higher dimensions, $k$-HodgeRank generalizes the comparison graph to a clique complex $\mathcal{K}_{n}(\mathcal{G})$ of dimension $k$ defined on $\mathcal{V}$. Instead of $\bm{\mathrm{s}}^{1}\in\mathbb{R}^{|\mathcal{E}|}$ assigning a real number to each edge ($1$-simplex), the data $\bm{\mathrm{s}}^{k}\in\mathbb{R}^{n_{k}}$ assigns a real number $s^{k}_{i}$ to each $k$-simplex $\sigma_{k}^{i}$ (a $(k+1)$-element subset of $\mathcal{V}$), where $i\in[n_{k}]:=\{0,\cdots,n_{k}-1\}$ and $n_{k}$ is the number of $k$-simplices in $\mathcal{K}_{n}(\mathcal{G})$. The solution to $k$-HodgeRank, the higher-dimensional analogue of Eq. (1), is
$\bm{\mathrm{s}}^{k-1}_{\ast}=\left(\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger}\right)^{+}\bm{\mathrm{B}}_{k}\bm{\mathrm{s}}^{k}.$ (2)
This vector assigns a score to each $(k-1)$-simplex in $\mathcal{K}_{n}(\mathcal{G})$. Similarly we define $\bm{\mathrm{s}}^{k}_{\mathrm{G}}:=\bm{\mathrm{B}}_{k}^{\dagger}\bm{\mathrm{s}}^{k-1}_{*}$, the projection of $\bm{\mathrm{s}}^{k}$ onto the space of conservative flows on the $k$-simplices of $\mathcal{K}_{n}(\mathcal{G})$. We define $k$-HodgeRank as finding $\bm{\mathrm{s}}^{k-1}_{*}$ given the following (classically stored) information: (i) $\bm{\mathrm{s}}^{k}$, (ii) a list of simplices in the clique complex $\mathcal{K}_{n}(\mathcal{G})$ (which includes the edges of $\mathcal{G}$).
Quantum least squares algorithms.—The HodgeRank problem in Eq. (1) is a type of least squares problem. For $k=1$, $k$-HodgeRank can be solved exactly with complexity $O(n^{3})$ [4]. Approximate algorithms, using graph Laplacian solvers [33, 34, 35] or topological signal processing (TSP) [21, 10], have complexities approximately quadratic in $n$, with varying dependence on the approximation parameters. For $k>1$, computing $\bm{\mathrm{s}}^{k}_{\mathrm{G}}$ and $\bm{\mathrm{s}}^{k-1}_{*}$ via least squares using existing classical sparse linear equation solvers has worst-case scaling $\Omega(n_{k})$ [36, 29], since $\bm{\mathrm{B}}_{k}$ is an $n_{k}\times n_{k-1}$ matrix with $(k+1)n_{k}$ non-zero entries. Generalizing the TSP approach similarly leads to algorithms with complexity $\Omega(n_{k})$ [37]. Since $n_{k}$ can grow like $\binom{n}{k+1}$, this can lead to exponential scaling in $k$.
There are several quantum algorithms for least square problems [27, 28, 38, 39, 40]. The best known algorithm requires $\tilde{O}(\alpha_{\bm{\mathrm{A}}}\kappa_{\bm{\mathrm{A}}}\log(1/\varepsilon))$ calls to a block encoding of $\bm{\mathrm{A}}$, where $\alpha_{\bm{\mathrm{A}}}$ is the scaling factor coming from the block encoding, $\kappa_{\bm{\mathrm{A}}}$ is (an upper bound on) the effective condition number of $\bm{\mathrm{A}}$, and $\varepsilon$ is the accuracy [40]. Their algorithm combines QSVT and variable-time amplitude amplification (VTAA) [41], modifying the previous approach [38].
These algorithms do not lead to efficient quantum algorithms for $k$-HodgeRank. Known block encodings of $\bm{\mathrm{B}}_{k}$ are constructed via quantum random access memory (QRAM) or a sparse access input model (SAIM). In these cases, $\alpha_{\bm{\mathrm{B}}_{k}}$ scales with either the Frobenius norm or the sparsity of $\bm{\mathrm{B}}_{k}$ [38, 40], which both have worst-case scaling $\Omega(n^{k})$ as discussed previously. We summarize these relationships between classical least squares algorithms, quantum least squares algorithms, HodgeRank, and quantum $k$-HodgeRank in Figure 1 in Supplemental Material [32].
We present a quantum algorithm to solve $k$-HodgeRank. Given a quantum state that encodes $\bm{\mathrm{s}}^{k}$ and a (classically stored) list of edges in the clique complex, our algorithm outputs a quantum state that approximates a state proportional to Eq. (2) using QSVT [42] and a projected unitary encoding (PUE) of $\bm{\mathrm{B}}_{k}$. Its complexity has quadratically worse dependence on the effective condition number than Ref. [40] but a simpler implementation. Moreover, the PUEs of $\bm{\mathrm{B}}_{k}$ and $\bm{\mathrm{B}}_{k+1}$ have scaling factors $\sqrt{n}$ [43] or $n$ [44] depending on implementation, thus avoiding the potentially exponential complexity arising from QRAM/SAIM. While the algorithm presented in Ref. [40] could possibly be adapted to work for PUEs while maintaining improved dependence on the effective condition number, we avoid this for simplicity.
Quantum HodgeRank.— We now give our quantum algorithm for $k$-HodgeRank (quantum $k$-HodgeRank) that scales exponentially better than its classical counterparts in $k$. It prepares a quantum state that encodes an approximate solution to Eq. (2) with a circuit depth that does not, in general, depend on the dimension $k$. Our algorithm is a classically controlled quantum circuit, as the structure is determined by $\mathcal{G}$.
We build quantum $k$-HodgeRank on the recently developed quantum algorithm for topological signal processing (QTSP) [45] for analyzing higher-order networks [19, 12, 13, 14, 15, 16]. We assume access to a quantum state preparation unitary $\bm{\mathrm{U}}_{\mathrm{prep}}$ that takes classical data $\bm{\mathrm{s}}^{k}$ and prepares
$\lvert\bm{\mathrm{s}}^{k}\rangle$ ::=:= $\frac{1}{\lVert\bm{\mathrm{s}}^{k}\rVert_{2}}\sum_{i\in[n_{k}]}\,s^{k}_{i}\>\lvert\sigma_{k}^{i}\rangle,$ (3)
where $\lvert\sigma_{k}^{i}\rangle$ is an $n$-qubit computational basis state representing $\sigma_{k}^{i}$ defined in [46, 44] (given in Eq. (18) in Supplemental Material [32]). We discuss the complexities of our applications in terms of the number of calls to $\bm{\mathrm{U}}_{\mathrm{prep}}$.
QTSP combines the quantum singular value transformation (QSVT) [42, 47] and PUEs $\bm{\mathrm{U}}_{\bm{\mathrm{\bm{\mathrm{B}}_{k}}}}$ and $\bm{\mathrm{U}}_{\bm{\mathrm{\bm{\mathrm{B}}_{k+1}}}}$ of the boundary matrices $\bm{\mathrm{B}}_{k}$ and $\bm{\mathrm{B}}_{k+1}$ [43, 48] (see Supplemental Material [32]). It prepares, upon successful postselection of the ancilla (which occurs with probability approximately $\mathcal{N}^{2}:=\big{\|}H\left(\bm{\mathrm{B}}_{k}/\sqrt{n},\bm{\mathrm{B}}_{k+1}^{\dagger}/\sqrt{n}\right)\lvert\bm{\mathrm{s}}^{k}\rangle\big{\|}_{2}^{2}$), the filtered state
$\lvert\bm{\mathrm{s}}^{k}_{\mathrm{fil}}\rangle:=\frac{1}{\mathcal{N}}\,H\left(\frac{\bm{\mathrm{B}}_{k}}{\sqrt{n}},\frac{\bm{\mathrm{B}}_{k+1}^{\dagger}}{\sqrt{n}}\right)\lvert\bm{\mathrm{s}}^{k}\rangle.$ (4)
Here $H(x,y):[-1,1]^{2}\rightarrow[-1,1]$ is a sum of two real polynomials $p(x)$ and $q(y)$, each with degree at most $D$, applied to the singular values of the respective matrices, where $p$ and $q$ are either both even or one is zero. The polynomial $H(x,y)$ is chosen so that |$\ket{\bm{\mathrm{s}}^{k}_{\mathrm{fil}}}$ extracts specific information from |$\ket{\bm{\mathrm{s}}^{k}}$. QTSP requires one call to $\bm{\mathrm{U}}_{\mathrm{prep}}$ as well as $O\big{(}D\big{)}$ calls to $\bm{\mathrm{U}}_{\bm{\mathrm{B}}_{k}}$, $\bm{\mathrm{U}}_{\bm{\mathrm{B}}_{k+1}}$, their inverses, and $O(D)$ applications of $\mathrm{C}_{\Pi_{j}}\mathrm{NOT}$ for $j\in\{k-1,k,k+1\}$, where $\Pi_{j}$ is the projector onto the states corresponding to $j$-simplices [49, 50], as well as post-selection on the correct ancilla state. The $\mathrm{C}_{\Pi_{j}}\mathrm{NOT}$ gates depend on the edges in the underlying graph $\mathcal{G}$ of the clique complex $\mathcal{K}_{n}(\mathcal{G})$ and are implemented via classical control. The circuit for QTSP has non-Clifford gate depth $O(Dn\log n)$ and uses $O(n)$ qubits. We review QTSP and its complexity in Supplemental Material [32] (see [45]).
Quantum $k$-HodgeRank utilizes QTSP to take as input the state |$\ket{\bm{\mathrm{s}}^{k}}$ as in Eq. (3) and output a quantum state |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{\ast}}$ that approximates
$\lvert\bm{\mathrm{s}}^{k-1}_{\ast}\rangle:=\frac{1}{\mathcal{N}_{\ast}}\,{{\sum}}_{i\in[n_{k-1}]}\,{s}^{k-1}_{\ast,i}\,\lvert\sigma_{k-1}^{i}\rangle$ (5)
where $\mathcal{N}_{\ast}:=\big{\|}\big{(}\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger}\big{)}^{+}\bm{\mathrm{B}}_{k}\lvert\bm{\mathrm{s}}^{k}\rangle\big{\|}_{2}$. We apply an approximation of $\big{(}\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger}\big{)}^{+}\bm{\mathrm{B}}_{k}$, given by a polynomial $H\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n},\bm{\mathrm{B}}_{k+1}^{\dagger}/\sqrt{n}\big{)}=p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}$ solely in $\bm{\mathrm{B}}_{k}$, to the input state. The polynomial $H(x,y)$ is defined by $q(y)=0$ and $p(x)=xg_{\varepsilon}(x^{2})$, where $\|g_{\varepsilon}(x)-1/(2\kappa_{k}^{2}x)\|\leq(\varepsilon/(2\kappa_{k}^{2}))$ for all $x\in[-1,-1/\kappa_{k}^{2}]\cup[1/\kappa_{k}^{2},1]$ (see [42, 51]). Here, $\kappa_{k}$ satisfies $\kappa_{k}^{-1}\in(0,\sqrt{n}/\xi_{\mathrm{min}}^{(k)})$, where $\xi_{\mathrm{min}}^{(k)}$ is the spectral gap (smallest nonzero singular value) of $\bm{\mathrm{B}}_{k}$, and $\varepsilon=(0,1/2)$. Importantly, $g_{\varepsilon}(x)$ is a polynomial of degree $O(\kappa_{k}^{2}\log(\sqrt{n}\kappa_{k}/\varepsilon))$, and
$\left\|\left(\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger}\right)^{+}\bm{\mathrm{B}}_{k}-\frac{2\kappa_{k}^{2}}{\sqrt{n}}\,p\left(\frac{\bm{\mathrm{B}}_{k}}{\sqrt{n}}\right)\right\|_{2}\leq\varepsilon.$ (6)
Let $\widetilde{\mathcal{N}}_{*}:=\big{\|}H(\bm{\mathrm{B}}_{k}/\sqrt{n})\ket{\bm{\mathrm{s}}^{k}}\big{\|}_{2}$. The output of quantum $k$-HodgeRank, after successful post-selection on the ancilla (which occurs with probability $\Omega\big{(}n\widetilde{\mathcal{N}}_{*}/(4\kappa_{k}^{4})\big{)}$), is
$\lvert\tilde{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle:=\frac{\sqrt{n}}{2\kappa_{k}^{2}\widetilde{\mathcal{N}}_{\ast}}\,{{\sum}}_{i\in[n_{k-1}]}\,\tilde{s}^{k-1}_{i,\ast}\,\lvert\sigma_{k-1}^{i}\rangle.$ (7)
This is a $2\varepsilon/\left(\mathcal{N}_{\ast}-\varepsilon\right)$-approximation to $\lvert\bm{\mathrm{s}}^{k-1}_{\ast}\rangle$ (in $\ell_{2}$ norm) if $\mathcal{N}_{*}>\varepsilon$. In cases where $\bm{\mathrm{s}}^{k}$ has high rankability (defined formally in Application 1), $\varepsilon$ can be chosen such that $\|\lvert\tilde{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle-\lvert{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle\|\leq 4\varepsilon/\mathcal{N}_{\ast}$ without significantly affecting the complexity (see Supplemental Material [32]). The complexity of QTSP implies that preparing |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$ from |$\ket{\bm{\mathrm{s}}^{k}}$, excluding post-selection, uses $O(n)$ qubits, a single call to $\bm{\mathrm{U}}_{\mathrm{prep}}$, and has non-Clifford gate depth
$O\big{(}n\kappa_{k}^{2}\log(n)\log(\sqrt{n}\kappa_{k}^{2}/\varepsilon)\big{)}.$ (8)
If $\kappa_{k}$ is small (e.g. not growing exponentially in $k$), this is much more efficient than classically computing $\bm{\mathrm{s}}^{k-1}_{*}$, even for small $k$. Examples of simplicial complexes with slowly-vanishing (or even constant in $k$) spectral gap are given in [44] and discussed in Supplemental Material [32]. Computing $p(\bm{\mathrm{B}}_{k})\bm{\mathrm{s}}^{k}$ classically requires $\tilde{O}(\kappa_{k}^{2})$ iterations of matrix-vector multiplication that each require $O(n_{k})$ operations [37]. This again highlights the dependency of the simplex dimension in the classical complexity, particularly for complexes with many $k$-simplices, i.e. when $n_{k}=\Theta\left(\binom{n}{k+1}\right)\approx\Theta(n^{k+1})$.
Similar to other quantum linear algebra algorithms [52, 38], the speedup of our algorithm comes from preparing the approximate solution state |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$. Extracting a precise classical representation of $\lvert\tilde{\bm{\mathrm{s}}}^{k-1}_{\ast}\rangle$ is generally expensive. We present applications that recover relevant information while avoiding full-state tomography where possible. These applications use quantum $k$-HodgeRank (or a straightforward modification) as a subroutine, and their complexities do not, in general, depend on $k$. This indicates that quantum $k$-HodgeRank, and QTSP in general, can potentially speed up specific computations defined on high-dimensional simplicial complexes. We assume that $\kappa_{k}$ and $\kappa_{k+1}$ of $\bm{\mathrm{B}}_{k}$ and $\bm{\mathrm{B}}_{k+1}$ are known and (ideally) small.
Approximating consistency measures.—We approximately measure the rankability of data in terms of so-called consistency and local inconsistency measures $\mathrm{R}(k)\in[0,1]$ and $\mathrm{R}_{\mathrm{C}}(k)\in[0,1]$, generalising definitions from [4]. The consistency measure associated to the data $\bm{\mathrm{s}}^{k}$ is the (Euclidean) length of $\bm{\mathrm{s}}^{k}_{\mathrm{G}}$ relative to $\bm{\mathrm{s}}^{k}$. This is analogous to a least squares “goodness of fit”. Similarly, the local inconsistency measure is the relative length of the projection of $\bm{\mathrm{s}}^{k}$ onto $\mathrm{im}(\bm{\mathrm{B}}_{k+1})$, which we denote $\bm{\mathrm{s}}^{k}_{\mathrm{C}}$. Explicitly,
$\mathrm{R}(k):=\left(\frac{\bm{\mathrm{s}}^{k}_{\mathrm{G}}\cdot\bm{\mathrm{s}}^{k}}{\|\bm{\mathrm{s}}^{k}\|^{2}}\right)^{1/2}\mbox{ and }\mathrm{R}_{\mathrm{C}}(k):=\left(\frac{\bm{\mathrm{s}}^{k}_{\mathrm{C}}\cdot\bm{\mathrm{s}}^{k}}{\|\bm{\mathrm{s}}^{k}\|^{2}}\right)^{1/2}.$ (9)
The value $\mathrm{R}(k)$ quantifies whether (sets of) alternatives can be meaningfully ranked based on the data, where larger values indicate high rankability. Conversely, $\mathrm{R}_{\mathrm{C}}(k)$ measures the presence of non-transitive preference relationships in the data, where the consensus prefers $A$ over $B$, $B$ over $C$, but $C$ over $A$. This has applications in finding arbitrage in currency exchange markets [4]. These measures are related to what is known as the Hodge decomposition (see Supplemental Material [32] for a brief introduction to the underlying topology).
We provide algorithms to compute $\varepsilon$-approximations to $\mathrm{R}(k)$ and $\mathrm{R}_{\mathrm{C}}(k)$ (see Supplemental Material [32] for details). Eq. (9) implies that $\mathrm{R}(k)=\left(\bra{\bm{\mathrm{s}}^{k}}\Pi_{\mathrm{G}}\ket{\bm{\mathrm{s}}^{k}}\right)^{1/2}$, where $\Pi_{\mathrm{G}}:=\bm{\mathrm{B}}_{k}^{\dagger}(\bm{\mathrm{B}}_{k}\bm{\mathrm{B}}_{k}^{\dagger})^{+}\bm{\mathrm{B}}_{k}$ projects onto $\mathrm{im}(\bm{\mathrm{B}}_{k}^{\dagger})$. Thus, we approximate $\mathrm{R}(k)$ using an accelerated Hadamard test [53] by using QTSP to build a PUE of (an approximate) $\Pi_{\mathrm{G}}$. Let $H(x,y)=p(x)=x^{2}g_{\varepsilon^{2}/2}(x^{2})$, where as before $g_{\varepsilon}$ is a polynomial which $\varepsilon/(2\kappa_{k}^{2})$-approximates $1/(2\kappa_{k}^{2}x)$. Then $\big{\|}\Pi_{\mathrm{G}}-2\kappa_{k}^{2}\,p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}\big{\|}_{2}\leq\varepsilon^{2}/2$, where $p(x)$ is degree $D=O\big{(}\kappa_{k}^{2}\log(n\kappa_{k}^{2}/\varepsilon^{2})\big{)}$ (see [45], Theorem 2). Thus, if $\bm{\mathrm{U}}_{\Pi_{\mathrm{G}}}=\mathrm{QTSP}(k,\mathcal{K}_{n},x^{2}g_{\varepsilon^{2}/2}(x^{2}))$ is the unitary that block encodes $p\big{(}\bm{\mathrm{B}}_{k}/\sqrt{n}\big{)}$ in the subspace defined by the ancilla in state |$\ket{0}^{\otimes a}$, then
$2\displaystyle 2\kappa_{k}^{2}\bra{\bm{\mathrm{s}}^{k}}\bra{0}^{\otimes a}\bm{\mathrm{U}}_{\Pi_{\mathrm{G}}}\ket{\bm{\mathrm{s}}^{k}}\ket{0}^{\otimes a}=\left(\mathrm{R}(k)\right)^{2}\pm\varepsilon^{2}/$ 2 . (10)
The value of $\mathrm{R}(k)$ can be $\varepsilon$-approximated with success probability $1-\delta$ using $O(\kappa_{k}^{2}\log(1/\delta)/\varepsilon^{2})$ controlled applications of $\bm{\mathrm{U}}_{\mathrm{prep}}$ and $\textsc{QTSP}\big{(}k,\mathcal{K}_{n},x^{2}g_{\varepsilon^{2}/2}(x)\big{)}$. Thus, if $\bm{\mathrm{U}}_{\mathrm{prep}}$ has non-Clifford gate depth $G$, the overall non-Clifford gate depth of this procedure is $O\left(\kappa_{k}^{2}\varepsilon^{-2}\log(1/\delta)\left(G+n\kappa_{k}^{2}\log n\log\left(\frac{n\kappa_{k}^{2}}{\varepsilon^{2}}\right)\right)\right)\approx\tilde{O}\left(\kappa_{k}^{2}\varepsilon^{-2}\log(1/\delta)\left(G+n\kappa_{k}^{2}\right)\right)$, where $\tilde{O}$ hides logarithmic factors of $n$, $\varepsilon$, and $\kappa_{k}$. The total runtime of the analogous approximate computation via classical TSP [21, 23] is $O\big{(}nkn_{k}\kappa_{k}^{2}\log(n\kappa_{k}^{2}/\varepsilon^{2})\big{)}$ [37, 10], where $n_{k}\leq\binom{n}{k+1}$.
This application can serve as a quantum-accelerated litmus test to check whether performing $k$-HodgeRank classically is worthwhile without the expensive computation required to determine $\mathrm{R}(k)$ exactly. For example, knowing $\mathrm{R}(k)$ to additive error $\varepsilon\approx 0.1$ would be indicative of whether the underlying data $\bm{\mathrm{s}}^{k}$ is amenable to a global ranking via $k$-HodgeRank.
Computing the local inconsistency measure is similar. In this task we estimate $\mathrm{R}_{\mathrm{C}}(k)=\left(\bra{\bm{\mathrm{s}}^{k}}\Pi_{\mathrm{C}}\ket{\bm{\mathrm{s}}^{k}}\right)^{1/2}$, where $\Pi_{\mathrm{C}}$ projects $\mathrm{im}(\bm{\mathrm{B}}_{k+1})$. Instead of applying $p(x)$ as defined above, apply $H(x,y)=q(y)=y^{2}g_{\varepsilon^{2}/2}(y^{2})$ such that $\|2\kappa_{k+1}^{2}\,q\big{(}\bm{\mathrm{B}}_{k+1}^{\dagger}/\sqrt{n}\big{)}-\Pi_{\mathrm{C}}\|\leq(\varepsilon^{2}/2)$. Analogously to approximating $\mathrm{R}(k)$, an accelerated Hadamard test can approximate $\mathrm{R}_{\mathrm{C}}(k)$ to additive error $\varepsilon$ with success probability $1-\delta$ using $\tilde{O}(\kappa_{k+1}^{2}\varepsilon^{-2}\log(1/\delta))$ applications of $\bm{\mathrm{U}}_{\mathrm{prep}}$ and $\textsc{QTSP}\big{(}k,\mathcal{K}_{n},y^{2}g_{\varepsilon^{2}/2}(y^{2})\big{)}$, with overall non-Clifford gate depth $\tilde{O}\left(\kappa_{k+1}^{2}\varepsilon^{-2}\log(1/\delta)\left(G+n\kappa_{k+1}^{2}\right)\right)$. The exact classical computation for this quantity, as given by Jiang et al. [4], is $O(n^{9})$ in the $k=1$ case, up to cubically worse in $n$. Implementing this polynomially classically has complexity $O\big{(}nkn_{k+1}\kappa_{k+1}^{2}\log(n\kappa_{k+1}^{2}/\varepsilon^{2})\big{)}$ for general $k$, with the same exponential scaling pitfalls as computing $\mathrm{R}(k)$ classically.
Relative rankings of few alternatives.—We can use quantum $k$-HodgeRank to approximate the scores of individual alternatives. This gives a heuristic for relatively ranking a small subset of alternatives without full-state tomography or computing classical $k$-HodgeRank. For more details see Supplemental Material [32].
Let $\bm{\mathrm{U}}_{k-1}^{i}$ be the tensor product of $\bm{\mathrm{X}}$ gates such that $\bm{\mathrm{U}}_{k-1}^{i}\ket{0\dots 0}=\ket{\sigma_{k-1}^{i}}$. By calling controlled applications of quantum $k$-HodgeRank, $\bm{\mathrm{U}}_{\mathrm{prep}}$, and $\bm{\mathrm{U}}_{k-1}^{i}$, a Hadamard test can approximate the score $s^{k-1}_{*,i}$ of each $(k-1)$-simplex $\sigma_{k-1}^{i}$. Approximating the scores for $L$ alternatives up to additive error $\varepsilon$ requires $O(n)$ qubits and $O\big{(}L\kappa_{k}^{2}\,\log(L/\delta)/(\sqrt{n}\varepsilon)\big{)}$ calls to each of $\textsc{QTSP}\big{(}k,\mathcal{K}_{n},xg_{\varepsilon/2}(x)\big{)}$, $\bm{\mathrm{U}}_{\mathrm{prep}}$, and $\bm{\mathrm{U}}_{k-1}^{i}$. Excluding the complexity of $\bm{\mathrm{U}}_{\mathrm{prep}}$, this circuit has non-Clifford gate depth is $\tilde{O}(L\kappa_{k}^{4}\sqrt{n}\log(L/\delta)/\varepsilon)$. One can relatively rank $L$ alternatives $\sigma_{k-1}^{i_{1}},\dots,\sigma_{k-1}^{i_{L}}\in\mathcal{K}_{n}(\mathcal{G})$ by approximating each score individually and ranking them according to these approximations.
This procedure is, in principle, efficient for small $L$. However, the number of $(k-1)$-simplices (i.e. basis states) is $O(n^{k})$, and so magnitude of the average amplitude may shrink exponentially in $k$. Thus, the accuracy and precision of $xg_{\varepsilon/2}(x^{2})$ and the Hadamard test may need to scale exponentially with $k$ to extract meaningful information from the approximate amplitudes, and a good choice of $\varepsilon$ cannot be determined a priori. While there may be specific choices of clique complex $\mathcal{K}_{n}(\mathcal{G})$ and data $\bm{\mathrm{s}}^{k}$ where the quantum advantage is greater, this is not generically the case.
Finding good alternatives.— We now discuss determining an approximate score for every $(k-1)$-simplex. As mentioned previously, full-state tomography is expensive. However, the tomography techniques given in [54] returns a classically stored $\ell_{\infty}$-norm approximation to |$\ket{\tilde{\bm{\mathrm{s}}}^{k-1}_{*}}$, denoted $\hat{\bm{\mathrm{s}}}^{k-1}_{*}$. As detailed in Supplemental Material [32], one can determine the score of every $(k-1)$-simplex, up to additive error $O(\varepsilon)$, in $O(\kappa_{k}^{4}n\varepsilon^{-2})$ applications of $\mathrm{QTSP}(k,\mathcal{K}_{n},xg_{\varepsilon}(x))$ and $\bm{\mathrm{U}}_{\mathrm{prep}}$. This $\ell_{\infty}$-norm approximation can be used to find alternatives with scores within additive error $O(\varepsilon)$ of the best alternative, or to determine whether the alternatives can be classified into subsets with similar scores within subsets and large differences in scores between subsets. However, like the previous application this has issues as the amplitudes of the state |$\ket{\bm{\mathrm{s}}^{k-1}_{*}}$ may be arbitrarily close, and an appropriate choice of $\varepsilon$ cannot be easily determined from |$\ket{\bm{\mathrm{s}}^{k}}$.
Dequantization. — Chia et al. [55] and Gharibian and Le Gall [56] have given different frameworks for dequantizing algorithms based on QSVT. In Supplemental Material [32] we show that these frameworks do not generically lead to efficient classical dequantizations of quantum $k$-HodgeRank or the computation of consistency or local inconsistency scores (Application 1). The reasons for this are: (i) difficulties obtaining polylogarithmic-time sampling and query access to $\bm{\mathrm{B}}_{k}$, and (ii) there exist simplicial complexes such that $\bm{\mathrm{B}}_{k}$ has Frobenius norm $\Theta(n^{k})$ and row sparsity $\Theta(n)$. We present a family of such complexes, previously studied by Berry et al. [44] in the context of quantum topological data analysis (QTDA). Each of these issues leads to inefficient dequantized algorithms in these frameworks. Methods for dequantizing QTDA rely on the relationship to estimating the dimension of the null space of an operator and do not obviously apply to our algorithms. An interesting future direction would be determining to what extent QTSP and quantum $k$-HodgeRank can be dequantized.
Conclusions. — Quantum $k$-HodgeRank is a straightforward combination of tools from QSVT and QTDA that approximates certain quantities much faster than classical algorithms for higher-order networks. HodgeRank offers an alternative solution to finding the stationary distribution of the random walk on the web page graph [4], and so quantum HodgeRank could be considered a variant of quantum PageRank algorithms [57, 58]. We note that while recommendation systems are related to rank aggregation problems, the problem formulation and methods used in HodgeRank differ significantly from those studied by Kerenidis and Prakash [39]. This distinction extends to the respective quantum algorithms.
Comparing to a broader range of classical algorithms is required to determine the presence of quantum advantage. Approximate methods for computing $\bm{\mathrm{s}}_{*}^{k-1}$, given in Eq. (2), exist for $k=1$ and some limited cases for $k=2$ [35, 36, 29]. However, these methods do not extend to Laplacian linear equations for $k\geq 2$, which are as hard to (approximately) solve as general linear equations [29], where sparse approximate solvers for linear equations with $N$ nonzero coefficients and condition number $\kappa$ run in time $O(\min\left\{N^{2.27159},N\kappa\right\})$ [29]. In the case of Laplacian equations, $N=\Omega(n_{k})$ can grow exponentially in $k$. This intuitively suggests that Equation (2) is hard to solve, even approximately.
The proof in [29] maps general linear equations to linear equations in the boundary matrix of a two-dimensional complex. This allows any least squares problem to be realised as a least squares problem in a boundary matrix $\bm{\mathrm{B}}_{2}$. Thus, theoretically quantum $2$-HodgeRank can solve general linear equations. Practically, the complexity of the mapping is linear in the number of variables and so this is inefficient. However, generalizing their methods to arbitrary $k$-dimensional complexes may offer a new avenue for efficient quantum linear systems solvers. A simpler intermediate step would be to characterize the boundary matrices of simplicial or clique complexes, in order to determine the class of least squares problems for which quantum $k$-HodgeRank could offer a speed-up compared to classical algorithms or QRAM/SAIM-based quantum methods.
Broadening the range of applications of discrete exterior calculus to higher-order networks would give quantum $k$-HodgeRank (and QTSP) more concrete problems where quantum $k$-HodgeRank is likely to perform best. Potential applications to the Kuramoto model of coupled oscillators [59] and higher-order analogues [19, 20], as well as variants of boundary value problems in fluid dynamics [60, 61, 62] and plasma physics [63, 64] on discretized and triangulated domains. Applications with quantum data would circumvent the difficulties in preparing the initial state |$\ket{\bm{\mathrm{s}}^{k}}$. Possible avenues include measuring quantum contextuality [65, 66] or studying multipartite interactions in complex quantum networks [67].
Acknowledgements
We thank Pedro C. S. Costa for several technical discussions. C.M.G.L. is supported by the Monash Graduate Scholarship and CSIRO Data 61 Top-Up Scholarship.